AThe city of Munich decides to urban an area for a new MERICAN ITshopping center. The costs of the land for the future parking places can be purchased for $400,000. The costs for the new roads are $560,000 and the annual maintenance expenses are $22,000. A new train station will cost $240,000. The annual operating costs for the train station are estimated to $175,000. KUWAIT The benefits are Tax from the shopping center $400,000/year Fee from the parking area $90,000/year Apply the B-C method by an estimated period of 10 years and MARR of 3%. Does it make sense? If The noises of the shopping center create disbenefits of $20,000, does this project still make sense?
Answer #1

We will try to evaluate BC ratio of above case and if BC ratio is greater than 1 then the project is feasible otherwise not.

We first calculate the present worth of benefits possible over the period of life of project

PW of Benefits from shopping center and parking area=490000(P/A,3%,10)=490000(8.530)

PW of Costs=Cost of land +Cost of new road+Cost for new train station+PW for maintenance+ PW for operating cost of train station

=400000+560000+240000+(175000+22000)(P/A,3%,10)=

=1200000+197000(8.530)

PW of Benefits-PW of Costs=293000(8.53)-1200000=1299290>0

Hence this project makes sense

Now if disbenefits is considered then PW of Costs will be increased by 20000(8.530)

B/C ratio=490000(8.53)/(1200000+217000(8.53))=1.369>1

Hence it makes project feasible in other case also